OpenStudy (anonymous):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below: f(x) = 72(1.25)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. (5 points) Part B: The table below shows the price f(t), in dollars, of product B after t years: t (# in years) 1 2 3 4 f(t) ($ in dollars) 65 84.5 109.85 142.81

OpenStudy (anonymous):

@nettle404

OpenStudy (anonymous):

@DeadReaper2073

OpenStudy (anonymous):

Dude, I am sorry but I suck at math. I failed it 2 times in a row,

OpenStudy (anonymous):

oh no. thanks anyway.

OpenStudy (anonymous):

@Spring98

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@Luigi0210 , @phi

OpenStudy (anonymous):

Could someone please help me?

OpenStudy (mathmate):

I suspect the function should be written: \(f(x) = 72(1.25)^x\) Whenever there is an exponent, write the ^ sign before it so helpers know that the number is exponentiated and not multiplied. It is perfectly ok to write f(x) = 72(1.25)^x which means the same function.

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

f(x) = 72(1.25)^x